# Video Unlearning via Low-Rank Refusal Vector

### v1

@@ 4.3 Qualitative Analysis
We report in Figure 2 a selection of qualitative results across the five considered unsafe categories:
@@
Figure 2: Qualitative results for the five unsafe categories. The top row shows uncensored video frames, while the bottom row shows corrected versions with our method.
4.4 Ablation Studies
We analyze the effect of varying two key parameters in our study: the rank of the cPCA operation, which determines the dimensionality of the subspace where the transformations are applied, and the $\lambda$ coefficient, which controls the strength of the unlearning intervention.
Figure 3: On the left, the figure illustrates the behavior of the Censorship Score as a function of the cPCA rank (see Eq. 4). On the right, the figure shows a decreasing trend in the Censorship Score as the value of $\lambda$ increases (Eq. 6).
cPCA rank.
The rank of the cPCA determines how strongly the information is compressed along the most relevant directions for the concept to be unlearned. Figure 3 (left) reports the censorship score as we vary the cPCA rank $k$. The score first decreases, reaching its minimum at $k=100$, and then increases again for lower ranks. When the cPCA rank $k$ is too large, higher than $100$, we also consider directions that are not or weakly connected to the target concept. Then the unwanted semantics become blended with other concepts, so the correction is no longer selective. Conversely, when the rank drops below $100$, we discard several principal directions along which the target concept is expressed, reducing correction effectiveness. An intermediate rank preserves all salient directions associated with the unwanted concept while excluding the ones associated with other semantics.
@@ 4.4 Ablation Studies
Impact of $\lambda$ on Censorship and Output Quality.
Figure 3 (right) shows that a larger suppression coefficient $\lambda$ lowers the censorship metric by attenuating the target unsafe concept. However, Figure 4 shows that the hyperparameter $\lambda$ also regulates a trade-off in output quality: moving left to right (increasing $\lambda$), small values leave much of the unsafe concept intact, whereas very large values over-suppress the latent code and visibly corrupt the video. We therefore adopt the intermediate setting $\lambda=1$, which provides the best compromise between effective censorship and visual fidelity.
5 Limitations
The ability to suppress specific concepts introduces risks of misuse, such as censoring legitimate content (e.g., historical documentation, artistic expression) or manipulating outputs to align with biased agendas. While our work focuses on technical safeguards, it does not address the ethical frameworks or governance structures needed to define and enforce “unwanted concepts”. Moreover, the directional nature of the refusal vector raises concerns about adversarial reversibility. Malicious actors could invert the vector to add erased concepts to the model or even implant new, harmful elements. Finally, our method removes one concept at a time.
@@ 6 Conclusions
Generative video diffusion models inherit risks from harmful content in their training data. We introduce the first training-free framework for targeted concept removal in video diffusion models. Our method utilizes a novel low-rank “refusal vector”, derived from minimal data (just 5 “safe”/“unsafe” image-prompt pairs) by analyzing latent differences in critical layers. This vector is embedded in the model weights to precisely suppress unwanted concepts like nudity, violence, or copyrighted material, without needing retraining or access to original data, and with no extra inference cost. Empirical evaluation demonstrates effective mitigation of harmful content while preserving overall video quality and prompt fidelity, highlighting the method’s precision and minimal impact.
Acknowledgements
We acknowledge support from Panasonic, the PNRR MUR project PE0000013-FAIR, and HPC resources provided by CINECA.
@@ Appendix A Technical Implementation Details
Open-Sora backbone. The diffusion backbone follows a Diffusion Transformer [30] topology with 19 double and 38 single residual blocks. A double block consists of two cross-attention sub-blocks, one for text prompt conditioning and the other for image conditioning, each laid out as: LayerNorm → Cross-Attention → LayerNorm → Feed-Forward Network (FFN). A single block contains one self-attention sub-block with the same LayerNorm → Self-Attention → LayerNorm → FFN pattern but no cross-modal conditioning.
Location of the weight injection. We apply the closed-form update of Eq. (8) to the FFN weights that immediately follow the two cross-attention modules in every double block:
@@
where $l\in\{17,18,19\}$ denotes the index of the double block (counting from the input side). No changes are applied to single blocks, as they hold self-attentive features rather than cross-modal semantics.
Why layers 17–19? A layer-wise sweep, injecting the correction into one double block at a time, shows that the most significant drop in unsafe content occurs when the edit is applied to the final three double blocks. Combining those layers (17-18-19) yields the best overall trade-off: censorship remains strong, yet visual quality is preserved. Extending the update to intermediate blocks (e.g., 14-19) keeps censorship high but introduces flicker and colour drift, and pushing the edit into earlier blocks degrades video quality monotonically. Similarly, adding the single blocks (layers beyond the 19) offers no benefit, so we update exactly the three final double blocks. This behaviour resonates with recent findings that high-level semantics emerge in the middle and late stages of deep architectures [25].
@@ Appendix B Model Weights update
In this section, we show that the subspace-aware edit of Eq. (6) can be absorbed into the parameters of the subsequent linear layer, detailing the steps to obtain the weights update in Eq. (7). We denote the pre-activation at layer $l+1$ in the original network as $\mathbf{x}^{l+1}=W^{l+1}\mathbf{x}^{l}$. Before this transform, we replace $\mathbf{x}^{l}$ with its corrected version $\tilde{\mathbf{x}}^{l}$ from Eq. (6). Substituting Eqs. (4)–(5) into (6) and propagating the result through $W^{l+1}$ yields the following sequence of equalities:
@@
=(6)${{=}}\mathbf{W}^{l+1}\!\left(\mathbf{x}^{l}\;-\;\lambda\Bigl{\langle}\mathbf{\hat{x}}^{\,l},\tfrac{\mathbf{\hat{r}}^{\,l}}{\lVert\mathbf{\hat{r}}^{\,l}\rVert_{2}}\Bigr{\rangle}\;\tfrac{\mathbf{r}^{*\,l}}{\lVert\mathbf{r}^{*\,l}\rVert_{2}}\right)$ ) (13)
[…]
@@
[…]
@@
=(6)${{=}}W^{\,l+1}\!\left(\mathbf{x}^{l}-\lambda\,\Bigl{\langle}\hat{\mathbf{x}}^{\,l},\tfrac{\hat{\mathbf{r}}}{\lVert\hat{\mathbf{r}}\rVert_{2}}\Bigr{\rangle}\tfrac{\mathbf{r}^{*}}{\lVert\mathbf{r}^{*}\rVert_{2}}\right)$ ) (21)
@@
Appendix C Additional Qualitative Results
@@
Appendix D Prompts for the refusal vectors

### v2

@@ 4.3 Qualitative Analysis
We report in Figure 2 a selection of qualitative results across the five considered unsafe categories on Open-Sora:
@@
Figure 2: Qualitative results for the five unsafe categories of T2VSafetyBench. Top: original uncensored frames. Bottom: corrected outputs with our method.
5 Discussion
In this section, we ablate the key components of our method. Specifically, we evaluate: (i) the impact of the cPCA subspace rank, (ii) the effect of the suppression strength controlled by $\lambda$, (iii) the sensitivity to the choice of safe/unsafe prompt pairs, (iv) the influence of neutral prompts used in cPCA, and (v) the effectiveness of cPCA compared to standard PCA in refining the refusal vector.
Effect of cPCA rank. The rank of the cPCA projection controls the dimensionality of the subspace in which unlearning is performed. Figure 3 (left) shows that censorship performance improves as the rank increases, reaching optimal results at $k=100$, before degrading again for higher or lower values. When $k$ is too large, unrelated directions are included, leading to entanglement with safe concepts. When $k$ is too small, relevant directions are discarded, reducing effectiveness. An intermediate rank ensures selective suppression of unsafe content while preserving safe semantics.
Impact of hyperparameters: layers, prompt pairs, $\lambda$. We apply refusal vectors to layers 17–18, which we found to be empirically the most effective positions for suppressing unsafe concepts without harming generation quality. Regarding the number of prompt pairs, we observe that increasing beyond five yields no noticeable performance; thus, we adopt five as the best trade-off between maintaining video quality and ensuring computational efficiency in deriving the refusal vector. Figure 3 (right) shows that a larger suppression coefficient $\lambda$ lowers the censorship metric by attenuating the target unsafe concept. However, Figure 4 also highlights the trade-off in output quality: small values leave unsafe content under-suppressed, while very large values over-suppress the latent code and visibly corrupt the video. We therefore adopt $\lambda=1$ as the best compromise between effective censorship and visual quality.
Figure 3: On the left, the figure illustrates the behavior of the censorship rate as a function of the cPCA rank (see Eq. 4). On the right, the figure shows a decreasing trend in the censorship rate as the value of $\lambda$ increases (Eq. 6).
@@ 5 Discussion
Sensitivity to Unsafe/Safe Prompt Pairs Choice. To assess whether performance depends on the semantic proximity of the five safe and unsafe pairs, we grouped prompts by cosine similarity in the Sentence-BERT (Reimers and Gurevych, 2019) embedding space: low (0.15–0.35), moderate (0.35–0.55), and high (0.55–0.75). Results in Table 5 show stable censorship scores over a representative subset of the “Pornography” category from T2VSafetyBench, with slightly better performance when prompts are highly similar, but consistently similar values across all the other ranges. This indicates that closer pairs provide a clearer estimate of the unsafe axis while reducing noise from unrelated content.
Stability to Neutral Set Choice in cPCA.
As with the safe/unsafe prompt pairs, our method is robust to the choice of neutral prompts in cPCA. Table 5 shows that across five random resamplings, the censorship rates vary by less than 2%, over a representative subset of the “Pornography” category from T2VSafetyBench. This indicates that the neutral set contributes to the contrastive subspace but does not require careful selection to achieve effective unlearning.
Refusal vectors vs. PCA and cPCA. Table 5 compares plain refusal vectors with PCA and cPCA refinements. While applying the refusal vector alone already reduces unsafe generations substantially (from 44.7% to 18.0% on the “Pornography” category), adding PCA provides an additional gain (16.9%). By contrast, cPCA achieves the strongest reduction (13.4%), showing that it better disentangles unsafe from safe semantics and avoids collateral suppression.
Table 3: Impact of semantic similarity between safe/unsafe pairs.
Similarity Censorship
Low (0.15–0.35) 25.3%
Moderate (0.35–0.55) 25.4%
High (0.55–0.75) 23.4%
Table 4: Stability of cPCA under different resamplings of the neutral set.
Resampling Censorship
cPCA 1 18.7%
cPCA 2 18.9%
cPCA 3 17.5%
cPCA 4 19.3%
cPCA 5 19.2%
Table 5: Comparison of refusal-only, PCA, and cPCA with respect to the baseline.
Method Censorship
Refusal Only 18.0%
+ PCA 16.9%
+ cPCA 13.4%
Baseline 44.7%
@@ 6 Conclusions
Generative video models inherit risks from harmful content in their training data. We introduce the first training-free closed-form weight update framework for targeted concept removal in video diffusion models. Our method utilizes a novel low-rank refusal vector, derived from a few “safe”/“unsafe” input pairs to derive a direction that encodes the target unwanted concept (e.g., nudity, violence, or copyrighted material). This vector is then embedded into the model weights as a closed-form update, without the need for retraining, access to original data, or any extra inference cost. Experiments on multiple established benchmarks show that the approach substantially reduces unsafe generations while preserving video quality and alignment with prompts.
7 Acknowledgement
We acknowledge partial financial support from Panasonic, the MUR FIS2 grant n. FIS-2023-00942 “NEXUS” (cup B53C25001030001), and the Sapienza grants RG123188B3EF6A80 (CENTS), RM1241910E01F571 (V3LI), and Seed of ERC grant “MINT.AI” (cup B83C25001040001). We acknowledge CINECA for computational resources and support. SF is co-funded by CINECA. MM is co-funded by ItalAI S.r.l.
Appendix
This document provides additional details to support the main paper. It includes the technical implementation details, the complete steps to derive the model weights update, extra qualitative results, and the prompt used to calculate the refusal vector.
We encourage readers to view the supplementary videos available in the index.html file inside the provided folder.
We additionally provide the code in the corresponding folder for reproducibility purposes.
@@ Appendix A Technical Implementation Details
Open-Sora backbone. The diffusion backbone of Open-Sora follows a Diffusion Transformer (Peebles and Xie, 2022) topology with 19 double and 38 single residual blocks. A double block consists of two cross-attention sub-blocks, one for text prompt conditioning and the other for image conditioning, each laid out as: LayerNorm → Cross-Attention → LayerNorm → Feed-Forward Network (FFN). A single block contains one self-attention sub-block with the same LayerNorm → Self-Attention → LayerNorm → FFN pattern but no cross-modal conditioning.
ZeroScopeT2V backbone. The diffusion backbone of ZeroScopeT2V follows the architecture of ModelScopeT2V (Wang et al., 2023b). The network is composed by a series of initial blocks, downsampling blocks, spatio-temporal blocks and upsampling blocks. Within blocks, cross-attention is employed to fuse information from conditioning modality, with a final FFN used to re-project embeddings.
Location of the weight injection. We apply the closed-form update to the weights of Eq. 8 the last FFN in the blocks for both the models. Particularly, in Open-Sora the intervention is performed over the cross-attention modules in every double blocks while in ZeroScopeT2V this operation is performed over up-blocks cross-attention.
@@
where $l$ denotes the index of the block considered. No changes are applied to single blocks, as they hold self-attentive features rather than cross-modal semantics.
@@ Appendix B Model Weights update
In this section, we show that the subspace-aware edit of Eq. 6 can be absorbed into the parameters of the subsequent linear layer, detailing the steps to obtain the weights update in Eq. 7. We denote the pre-activation at layer $l+1$ in the original network as $\mathbf{x}^{l+1}=W^{l+1}\mathbf{x}^{l}$. Before this transform, we replace $\mathbf{x}^{l}$ with its corrected version $\tilde{\mathbf{x}}^{l}$ from Eq. 6. Substituting Eqs. 45 into 6 and propagating the result through $W^{l+1}$ yields the following sequence of equalities:
@@
$=\mathbf{W}^{l+1}\!\left(\mathbf{x}^{l}\;-\;\lambda\Bigl\langle\mathbf{\hat{x}}^{\,l},\tfrac{\mathbf{\hat{r}}^{\,l}}{\lVert\mathbf{\hat{r}}^{\,l}\rVert_{2}}\Bigr\rangle\;\tfrac{\mathbf{r}^{*\,l}}{\lVert\mathbf{r}^{*\,l}\rVert_{2}}\right)$ (13)
$=\mathbf{W}^{l+1}\!\left(\mathbf{x}^{l}\;-\;\lambda\Bigl\langle U_{k}^{\top}\mathbf{x}^{l},\tfrac{\mathbf{\hat{r}}^{\,l}}{\lVert\mathbf{\hat{r}}^{\,l}\rVert_{2}}\Bigr\rangle\;\tfrac{U_{k}\mathbf{\hat{r}}^{\,l}}{\lVert U_{k}\mathbf{\hat{r}}^{\,l}\rVert_{2}}\right)$ (14)
@@
$=\mathbf{W}^{l+1}\!\left(\mathbf{x}^{l}\;-\;\lambda\,U_{k}\,\frac{\mathbf{\hat{r}}^{\,l}\mathbf{\hat{r}}^{\,l\!\top}}{\lVert\mathbf{\hat{r}}^{\,l}\rVert_{2}^{\,2}}\,U_{k}^{\top}\mathbf{x}^{l}\right)$ (16)
$=\underbrace{\mathbf{W}^{l+1}\Bigl(I-\lambda\,U_{k}\,\tfrac{\mathbf{\hat{r}}^{\,l}\mathbf{\hat{r}}^{\,l\!\top}}{\lVert\mathbf{\hat{r}}^{\,l}\rVert_{2}^{\,2}}\,U_{k}^{\top}\Bigr)}_{\displaystyle\tilde{\mathbf{W}}^{\,l+1}}\,\mathbf{x}^{l}$ (17)
@@
[…]
@@
Appendix C Additional Quantitative Results
We report additional quantitative results.
C.1 FVMD
Table 6 shows FVMD (Liu et al., 2024a) on T2VSafetyBench (lower is better). FVMD measures frame-to-frame consistency and motion stability via key-point tracking, addressing gaps in standard metrics that overlook temporal coherence. Sensitivity tests with injected noise and disrupted temporal order support its validity as a motion-consistency metric. Our method matches the Open-Sora baseline across all five categories, indicating preserved motion stability and temporal coherence.
Table 6: FVMD over the different target categories of T2VSafetyBench for the Open-Sora baseline.
Model Pornography Copyright Gore Sequential Action Risk Public figures
Baseline 14884.19 11600.61 13707.37 12665.67 11755.99
Censored 10527.19 7407.82 14690.15 12427.61 11048.42
C.2 Computational and Data Efficiency Comparison
Table 7 compares the computational requirements, estimated costs, and data needs of different approaches for concept removal and model adaptation. While full training and fine-tuning demand substantial GPU resources and large-scale datasets, our training-free method requires only a few prompt pairs and negligible compute, enabling fast and flexible updates.
Table 7: Comparison of methods. Abbreviations: FT = fine-tuning, TF = training-free, Data Req. = data requirements.
Method Description GPU h. Est. Cost Data Req. Flexibility
Open-Sora Training 3-stage full training $\sim$100k $\sim$\$200k 70M / 10M / 5M videos Not scalable
Open-Sora FT Targeted adaptation (25k clips) $\sim$1k $\sim$\$2k–4k 25k labeled videos Per-concept re-tuning
Ours (TF) Refusal vector + cPCA 0.4 $\sim$\$1 (10 fwd passes) 5 safe/unsafe pairs Fast, no data
C.3 Mass Erasure and Multi-vector unlearning
We assess whether multiple unsafe concepts can be censored simultaneously by composing per-layer projections along several refusal directions. Given concepts $\{c_{k}\}_{k=1}^{K}$ with corresponding refusal vectors $\{\mathbf{r}_{k}^{l}\}$ (one per layer $l$), we extend our update
$\tilde{\mathbf{x}}^{\,l}=\mathbf{x}^{\,l}-\lambda\Big\langle\mathbf{x}^{\,l},\frac{\mathbf{r}^{\,l}}{\|\mathbf{r}^{\,l}\|}\Big\rangle\frac{\mathbf{r}^{\,l}}{\|\mathbf{r}^{\,l}\|}$
to the multi-concept case as
$\tilde{\mathbf{x}}^{\,l}=\mathbf{x}^{\,l}-\sum_{k=1}^{K}\lambda_{k}\Big\langle\mathbf{x}^{\,l},\frac{\mathbf{r}_{k}^{\,l}}{\|\mathbf{r}_{k}^{\,l}\|}\Big\rangle\frac{\mathbf{r}_{k}^{\,l}}{\|\mathbf{r}_{k}^{\,l}\|}\,.$
We evaluate multiple categories both individually and in combination over a subset of 100 prompts, applying refusal vectors sequentially. With 1–4 refusal vectors, censorship improves substantially (e.g., Pornography: 63.0% to 26.0%; Copyright: 69.0% to 32.0%; Public Figures: 33.0% to 7.0%), while FVD increases only moderately. Using five vectors leads to more aggressive edits and a sharper FVD rise, particularly on action-risk prompts where temporal dynamics are harder to preserve. Overall, higher-fidelity categories remain more robust than structurally complex ones, indicating that the method interacts consistently across semantic domains. Table 8 reports the full results.
Table 8: Censorship, FVD, and $M$-Notox under sequential refusal-vector combinations.
Pornography Copyright Public Figures Sequential Action Risk Gore
Method Cens. (%) FVD $M$-Notox Cens. (%) FVD $M$-Notox Cens. (%) FVD $M$-Notox Cens. (%) FVD $M$-Notox Cens. (%) FVD $M$-Notox
1 Refusal vector 26.0 191.92 20.46 41.0 214.72 20.77 1.0 233.66 20.39 14.5 273.40 20.58 13.0 230.30 21.10
2 Refusal vectors 27.0 195.49 21.56 42.0 213.59 22.20
3 Refusal vectors 27.0 192.59 21.44 41.0 212.82 21.98 8.0 216.92 21.16
4 Refusal vectors 26.0 211.46 21.30 32.0 226.88 22.01 7.0 266.38 21.19 14.5 250.29 21.40
5 Refusal vectors 16.0 209.86 21.37 8.0 245.13 21.29 2.0 239.74 20.84 10.9 579.37 21.14 15.0 254.31 20.76
OpenSora 63.0 183.44 20.49 69.0 201.76 22.03 33.0 212.89 21.64 45.5 220.54 22.32 87.0 236.36 21.67
Appendix D Additional Qualitative Results
@@
Appendix E Prompts for the refusal vectors
@@
Appendix F Applicability of UCE to T2V
Methods such as UCE ((Gandikota et al., 2024)) are originally designed for text-to-image pipelines and rely on architectural assumptions that do not transfer directly to text-to-video models (e.g., single-frame UNets, CLIP-based spatial cross-attention, and the absence of temporal components). Since approaches like UCE act through text-conditioned concept directions, they can nonetheless be adapted to a video setting. To examine this possibility, we implemented the closest feasible adaptation within a T2V pipeline: the UCE edit is applied to the text encoder, and the modified representations are propagated through the full video generator. We report the results on the Pornography category of T2VSafetyBench in Table 9. The retain-prompt set we used is: {woman; man; person; face; smile; body; lady; girl; shows; eyes; nose; look; move}, for a total of 13 preserve prompts. Following the official implementation, we used $\texttt{erase\_scale}=1$.
Table 9: Evaluation of UCE adapted to text-to-video. Comparison on the Pornography category of T2VSafetyBench between UCE and our proposed method.
Censorship $\downarrow$ FVD $\downarrow$ $M$-Notox $\downarrow$
Category ZeroScopeT2V UCE Ours ZeroScopeT2V UCE Ours ZeroScopeT2V UCE Ours
Sexual 51.5% 8.9% 9.1% 60.96 103.3 63.51 22.26 23.43 22.62
UCE achieves a censorship rate comparable to our method, but at a substantially higher cost in video quality: its FVD rises from 63.51 to 103.3, corresponding to an increase of more than 60%. In addition, the MM-Notox score is higher, indicating greater semantic drift from the intended prompt. These results are consistent with the architectural constraints of UCE’s original formulation and highlight the benefits of a T2V-specific approach.
F.1 UCE Hyperparameter Analysis
For completeness, we report the full details of the UCE adaptation and the accompanying hyperparameter study regarding Table 9. Results are shown in Table 10.
The chosen configuration in Table 9 used a retain-prompt set consisting of 13 prompts, a $\texttt{erase\_scale}=1$ and $\texttt{preserve\_scale}=1$. We explored $\texttt{preserve\_scale}\in\{0.5,1,5,30\}$ (denoted $\lambda_{p}$ in Table 9), testing each value with both our 13-prompts configuration and an additional 60-prompts retain set.
With 13 preserve prompts, larger $\lambda_{p}$ values modestly improves censorship but cause video quality to deteriorate sharply: FVD rises from 103 at $\lambda_{p}=1$ to over 300 for $\lambda_{p}\in\{5,30\}$, more than three times worse than our released configuration. With 60 preserve prompts, smaller $\lambda_{p}$ values lead to full suppression, but the resulting videos exhibit very poor quality (FVD $\approx 470$490490). This suggests that expanding the preserve set forces the model into an over-suppressive regime that achieves safety only at the cost of drastic fidelity loss.
Across conditions, the only configuration approaching our censorship level is the combination (60 prompts, $\lambda_{p}=30$), which achieves 5.9% censorship but remains substantially inferior in video quality (FVD =282.59=282.59 vs. $63.51$). Thus, among all tested hyperparameters, the chosen configuration (13 prompts, $\lambda_{p}=1$) consistently provides the best balance of censorship, fidelity, and semantic stability.
Table 10: UCE hyperparameter study. Censorship, video quality (FVD), and semantic drift ($M$-Notox) across different values of $\lambda_{p}$ and preserve–prompt set sizes (13 vs. 60). The configuration used in 9 corresponds to 13 preserve prompts and $\lambda_{p}=1$, which achieves the best trade-off between suppression, fidelity, and semantic stability.
$\lambda_{p}$ # Preserve Prompts Cens. (%) FVD $M$-Notox
0.5 13 0 463.06 23.50
1 13 8.9 103.30 23.43
5 13 1.2 307.98 24.86
30 13 2.4 309.95 24.13
0.5 60 0 469.64 24.46
1 60 0 486.80 23.65
5 60 2.4 407.17 23.67
30 60 5.9 282.59 20.14
Ours 9.1 63.51 22.62
ZeroScopeT2V 51.5 60.69 22.26
Appendix G Neutral Concept Selection in cPCA
A potential concern in contrastive PCA is the selection of “neutral concepts,” particularly when semantic overlap may occur with the unsafe concepts of interest (e.g., knife vs. violence). In such cases, subtracting the neutral covariance $C_{e}$ could unintentionally weaken the unsafe direction, thereby reducing suppression effectiveness. In our approach, approximately 1000 neutral prompts are generated using GPT-based sampling. The sampling process explicitly requests a broad set of generic and unrelated concepts, ensuring a diverse and domain-agnostic neutral space without requiring manual curation. To examine the robustness of this procedure, we conduct multiple resamplings of the neutral prompt set (see Table 5). Across repetitions, censorship performance varies minimally, indicating that neutrality selection does not significantly influence results. We also analyze the impact of the contrast parameter $\alpha$ on OpenSora over Pornography category. As shown below, performance remains stable across a practical range of $\alpha$ values, confirming that the method is resilient to configuration choices in neutral concept selection.
Table 11: Effects on censorship rate for different alpha values for cPCA
$\alpha$ Censorship rate (%)
0.5 15.1
1.0 13.4
1.5 14.7
Appendix H Robustness to Adversarial Attacks
Adversarial Prompting Evaluation.
Adversarial prompting represents an important dimension of safety for text-to-video systems. The benchmarks employed in our study (T2VSafetyBench and SafeSora) already incorporate adversarial-like inputs generated through diverse mechanisms. In particular, Miao et al. (2024) constructs its dataset from 3 sources: real user NSFW prompts; GPT-4–generated malicious prompts; and three different jailbreak attacks targeting diffusion models. SafeSora likewise includes complex unsafe descriptions involving multi-element scenes and narrative contexts, extending well beyond isolated keyword triggers.
To further assess robustness, we applied the adversarial attack from Zhang et al. (2024b) to a subset of 100 Pornography prompts for which our method produced 0% unsafe frames under the GPT-4o judge, i.e. prompts on which the refusal direction achieves complete censorship prior to any attack. After the adversarial attack, the unsafe-frame rate rose to only 4%. This small increase indicates that the training-free unlearning remains effective and that the learned refusal direction retains substantial stability under targeted prompt optimization.